# Landau-de Gennes Modelling of Confinement Effects and Cybotactic Clusters in Bent-Core Nematic Liquid Crystals

### v1

@@ Abstract
We study bent-core nematic (BCN) systems in two-dimensional (2D) and three-dimensional (3D) settings, focusing on the role of cybotactic clusters, phase transitions, confinement effects and applied external fields. We propose a generalised version of Madhusudhana’s two-state model for BCNs in [1] with two order parameters: $\mathbf{Q}_{g}$ to describe the ambient ground-state (GS) molecules and $\mathbf{Q}_{c}$ to describe the ordering within the cybotactic clusters. The equilibria are modelled by minimisers of an appropriately defined free energy, with an empirical coupling term between $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$. We demonstrate two phase transitions in spatially homogeneous 3D BCN systems at fixed temperatures: a first-order nematic-paranematic transition followed by a paranematic-isotropic phase transition driven by the GS-cluster coupling. We also numerically compute and give heuristic insights into solution landscapes of confined BCN systems on 2D square domains, tailored by the GS-cluster coupling, temperature and external fields. This benchmark example illustrates the potential of this generalised model to capture tunable director profiles, cluster properties and macroscopic biaxiality.
@@ I Introduction
BCNs were discovered in the 1990’s; they are remarkable materials that can exhibit different types of order — polar, chiral, octupolar, biaxial order etc. They exhibit exotic twist-bend, splay-bend, smectic and columnar phases, amongst many others [3]. These BCN phases offer tremendous potential for electro-optic devices, rheological applications and novel functional materials. Cybotactic clusters are an intrinsic feature of BCNs; they are smectic nanoclusters such that BCN molecules get locked into clusters because of their kinked shape. The existence of cybotactic clusters has been confirmed by multiple experimental methods, e.g., SAXS, cryto-TEM methods, electro-optic measurements [6, 7, 8] etc. Cybotactic clusters are small, usually tens of nanometers in lateral dimensions and persist in the nematic and sometimes, even in the isotropic BCN phase. Cybotactic clusters render additional ordering to the conventional nematic phase and this ordering clearly manifests in optical images (four-lobed structures in SAXS measurements), rheological properties (non-Newtonian properties), elastic properties (suppressed twist elastic constants leading to spontaneous chirality), macroscopic biaxiality [3] etc. As such, it is essential to quantify the additional ordering induced by the cybotactic clusters and how this ordering modulates the conventional ambient nematic ordering.
@@
In [1], Madhusudhana proposes a two-state model for BCNs, based on the hypothesis that there are two types of molecules: the ground-state (GS) molecules that constitute the out-of-cluster ambient environment and the excited molecules (ES) that constitute the cybotactic clusters. There are two scalar order parameters, $S_{g}$ and $S_{c}$, to describe the orientational ordering of the GS molecules and the ES molecules respectively. Madhusudhana assumes spatially uniform $S_{g}$ and $S_{c}$ profiles, i.e., he works with spatially homogeneous BCN systems along with various other assumptions — fixed time-independent and temperature-independent cluster sizes, an empirical coupling term between $S_{g}$ and $S_{c}$ and no explicit information about the BCN directors. Madhusudhana proposes a phenomenological free energy comprising two Ginzburg-Landau type potentials for $S_{g}$ and $S_{c}$ with various temperature-dependent and material-dependent parameters and an empirical GS-ES coupling term that is parameterised by a coupling coefficient. The physically observable spatially homogeneous BCN configurations are modelled by minimisers of this two-state free energy. The most important finding concerns the existence of a new paranematic phase with small values of $S_{g}$ and $S_{c}$, just before the transition to the disordered isotropic phase. This model also effectively captures other experimental trends — enhanced order parameters induced by the cybotactic clusters, a nematic-paranemetic phase transition followed by a paranematic-isotropic phase transition with increasing temperature and non-linear dependence of order parameters on external magnetic fields.
@@
However, Madhusudhana’s model cannot account for confinement effects or spatial inhomogeneities. In [9], the authors propose a one-dimensional model for spatially inhomogeneous BCN samples, including elastic energy terms for $S_{g}$ and $S_{c}$. They study BCNs in channel geometries with fixed boundary conditions for $S_{g}$ and $S_{c}$; their numerical results demonstrate that confinement or geometric frustration promotes interior clustering or enhances interior ordering and that the GS-ES coupling can promote ordered director profiles even at high temperatures contrary to the expectation of isotropy or disorder at high temperatures. In [10], the authors study the effects of doping on cybotactic clusters along with the emergence of polar order in such BCN systems. They perform extensive experiments on phase transitions, birefringence, order parameter measurements and dielectric spectroscopy data and propose a simple Madhusudhana-type free energy to partially describe their experimental results. In this paper, we build on these works to propose a generalised free energy for confined BCN systems in two-dimensional (2D) and three-dimensional (3D) settings. By analogy with Landau-de Gennes theory, we define two tensor order parameters: $\mathbf{Q}_{g}$ to describe the orientational ordering of the GS molecules and $\mathbf{Q}_{c}$ to describe the orientational ordering of ES molecules within cybotactic clusters [11]. The key advantage is that $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ not only contain information about $S_{g}$ and $S_{c}$ but also contain information about the GS and ES directors, that are the preferred directions of averaged alignment of the GS and ES molecules (within clusters) respectively. The free energy has an elastic energy to penalise spatial inhomogeneities and a bulk energy that determines the GS and ES ordering as a function of temperature, GS-ES coupling and other parameters, along with terms that account for external (electric or magnetic) fields. The bulk energy is a direct generalisation of Madhusudhana’s free energy and is composed of Ginzburg-Landau type potentials for $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ respectively and a $\mathbf{Q}_{g}$-$\mathbf{Q}_{c}$ coupling term. The Ginzburg-Landau potentials dictate that $S_{g}$ and $S_{c}$ increase with decreasing temperatures and approach zero with increasing temperature (the isotropic or disordered phase). The GS-ES coupling and external field terms can strongly boost ordering, even for high temperatures. The physically observable textures are modelled by minimisers of this generalised energy, subject to the imposed boundary conditions. The minimisers and the critical points of the free energy are solutions of boundary-value problems for the associated Euler-Lagrange equations — typically a system of coupled and nonlinear partial differential equations and the rest of this manuscript is devoted to the study of critical points of this generalised free energy in physically motivated settings.
@@
The rest of the paper is organised as follows. In Section II, we introduce a generalised Landau-de Gennes type free energy for BCN systems. In Section III, we study phase transitions in spatially homogeneous BCN samples, with and without external fields, in 2D and 3D. Notably, we find two phase transitions in spatially homogeneous 3D BCN samples driven by increasing GS-ES coupling strength, at high temperatures. The GS-ES coupling strength is measured in terms of a parameter $\gamma$. If $\gamma=0$, then $S_{g}=S_{c}=0$ for spatially homogeneous samples at high temperatures (defined by $A>0$ in 5). As $\gamma$ increases, the energy minimisers have positive values of $S_{g}$ and $S_{c}$ even for high temperatures. At a fixed high temperature, increasing $\gamma$ has a strong ordering effect that counteracts the disordering effects of high temperature. This is analogous to the nematic-paranematic phase transition and the paranematic-isotropic phase transition with increasing temperature, at fixed $\gamma$, as reported in [1]. The external fields typically boost the values of $S_{g}$ and $S_{c}$, shift the bifurcation points and we lose the paranematic-isotropic phase transition. Given that the values of $S_{g}$ and $S_{c}$ are very small in the paranematic phase, it is unclear if the paranematic phase would be distinguishable from the conventional isotropic phase in experiments. In Section IV, we study solution landscapes of confined BCN systems on a square domain with Dirichlet boundary conditions for $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$. This is a benchmark well-studied example for conventional nematics; see the papers [12, 13, 14]. A solution landscape describes a connected family of critical points of the generalised free energy for this model problem, tracing the pathways from the unstable saddle points to the stable energy minimisers. Of course, the unstable saddle points can play a crucial role in the system dynamics and the selection of the energy minimiser. The GS-ES coupling term can stabilise symmetric structures and new ordered profiles at high temperatures. We study the interplay between domain size, temperature, the GS-ES coupling and in some cases, external field strength on the solution landscapes. We recover some familiar $\mathbf{Q}_{g}$ profiles from the studies on conventional nematics, some new $\mathbf{Q}_{g}$ profiles tailored by the additional physics in this model, co-aligned $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ profiles with co-aligned GS and ES directors and antagonistic $(\mathbf{Q}_{g},\mathbf{Q}_{c})$-profiles with mutually perpendicular GS and ES directors. Our simple examples illustrate the tremendous potential of tuning model parameters to get exotic morphologies, some of which could lead to macroscopic biaxiality e.g., the antagonistic $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ profiles. We conclude with a discussion of the limitations of the model and future improvements in Section VI.
@@ II Landau-de Gennes Modelling for Bent-Core Nematic Systems
In this paper, we propose a two-state generalised Landau-de Gennes (LdG) type free energy for confined BCN systems building on Madhusudhana’s model in [1]. We define a GS LdG order parameter, $\mathbf{Q}_{g}$, which is a symmetric, traceless $d\times d$ matrix, where $d$ is the spatial dimension. The eigenvectors of $\mathbf{Q}_{g}$ model the GS nematic directors, interpreted as the locally preferred directions of averaged GS molecular alignment in space and the associated eigenvalues measure the degree of orientational ordering about the GS directors. We define the GS director to the eigenvector of $\mathbf{Q}_{g}$ with the largest positive eigenvalue. The GS is said to be in the isotropic phase if $\mathbf{Q}_{g}=0$. It is in the uniaxial phase if $\mathbf{Q}_{g}$ has two equal non-zero eigenvalues for $d=3$ and $\mathbf{Q}_{g}$ can be written as
@@
Analogously, we define an ES LdG order parameter, $\mathbf{Q}_{c}$, to describe the state of ordering of the ES molecules and based on the assumption that the ES molecules form cybotactic clusters, we refer to $\mathbf{Q}_{c}$ as the cluster order parameter. For example, in $d=2$, we have
@@
where $\mathbf{n}_{c}$ is the ES/cluster director and $S_{c}$ is positive. As an immediate consequence, we have two tensor order parameters, $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$, as opposed to two scalar order parameters, $S_{g}$ and $S_{c}$ in [1], that can describe complex information about the GS and ES directors, structural transitions and confinement effects in inhomogeneous systems which are outside the scope of the two-state model in [1].
@@
where $\Omega$ is a $d$-dimensional bounded domain with Lipschitz boundary and $d=2,3$. The first two terms are the elastic energy density terms associated with $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ respectively, i.e., we assume a one-constant elastic energy density to penalise spatial inhomogeneities in $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ such as strong deformations of the clusters, interior defects etc. We note that other choices of the elastic energy density are also possible, see [15]. $K_{g}$ and $K_{c}$ denote the positive elastic constants of the GS and ES molecules, respectively. The mole fraction of the ES molecules $a_{x}$ is given by
@@
as $E_{ex}$ is the excitation energy of the ES molecules, $k_{B}$ is Boltzmann’s constant and $T$ is the temperature. $a_{x}=0.1$ throughout the paper consistent with [9]. The parameters $a_{g}$, $B_{g}$, $C_{g}$, and $T^{*}$ are the usual LdG parameters that describe the first-order nematic-isotropic transition for the GS molecules and $\gamma$ is the coupling parameter between the GS molecules and the clusters [9, 1]. We note that the term $\textrm{tr}\left(\mathbf{Q}_{g}\mathbf{Q}_{c}\right)$ is a generic coupling term for systems with two order parameters that is consistent with frame-indifference and material symmetry requirements [16]. We are unable to comment on whether this coupling accurately captures the physics of the GS-ES interactions. $\alpha_{c}$ and $\beta_{c}$ are coefficients for the saturation terms to ensure that $|\mathbf{Q}_{c}|^{2}$ is reasonably bounded, for at least a certain range of $\gamma$. In the subsequent numerical simulations, we adjust their values so that the resulting values of $S_{c}$ are reasonable and not too large. In each cluster, $N_{c}$ represents the number of ES molecules. We also consider the effects of an external field, $\mathbf{E}=E_{a}\mathbf{e}$, whose direction is modelled by a unit-vector $\mathbf{e}$, and the magnitude is $E_{a}$. $E_{el}$ is the external field energy density $(\frac{1}{2}\epsilon_{0}\Delta\epsilon E_{a}^{2})$ where $\epsilon_{0}$ is the free-space permittivity, $\Delta\epsilon$ is a material anisotropy parameter (dielectric or magnetic) for GS molecules and $J$ accounts for the larger shape anisotropy of ES molecules, i.e., $J>1$. In particular, the GS molecules tend to align with $\mathbf{E}$ if $\Delta\epsilon>0$ and orthogonal to $\mathbf{E}$ if $\Delta\epsilon<0$ [2].
@@
and $\tilde{\Omega}$ is the rescaled domain. For comparison with previous work on one-dimensional BCN systems in [9], we set $A=\pm 0.04$, $B=0.34$, $C=0.9$, $D=2\times 10^{-3}$, $E=1.6\times 10^{-4}$, $M=9.7\times 10^{-5}$, $N=1.78\times 10^{-3}$, $P=0.133$ which are derived from $K_{g}=K_{c}=K=15pN=15\times 10^{-12}N=15\times 10^{-7}\text{dyn}\text{(under one constant approximation)}$; $a_{g}=0.04\times 10^{7}/4\ cgs$, $B_{g}=1.7\times 10^{7}/4\ cgs$, $C_{g}=4.5\times 10^{7}/4\ cgs$, $\alpha_{c}=0.22\times 10^{7}/4\ cgs$, $\beta_{c}=4.0\times 10^{7}/4\ cgs$, and $\gamma=5.0\times 10^{7}/4\ cgs$; $T^{*}=355K$, $N_{c}=50$, $J=1.2$, $E_{el}=2000ergs/cm^{3}$, $E_{ex}=1.1\times 10^{-13}ergs$, and $T=360K$ (for the high temperature case with $A>0$ ) and $350K$ (for the low temperature case with $A<0$). In subsequent sections, we focus on phase transitions in spatially homogeneous systems and confinement effects in square domains as a function of temperature $A$, the GS-ES coupling strength measured by $D$ and external field strength, $E$.

### v2

@@ Abstract
We study bent-core nematic (BCN) systems in two-dimensional (2D) and three-dimensional (3D) settings, focusing on the role of cybotactic clusters, phase transitions, confinement effects and applied external fields. We propose a generalised version of Madhusudana’s two-state model for BCNs in [17] with two order parameters: $\mathbf{Q}_{g}$ to describe the ambient ground-state (GS) molecules and $\mathbf{Q}_{c}$ to describe the additional ordering induced by the cybotactic clusters. The equilibria are modelled by minimisers of an appropriately defined free energy, with an empirical coupling term between $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$. We demonstrate two phase transitions in spatially homogeneous 3D BCN systems at fixed temperatures: a first-order nematic-paranematic transition followed by a paranematic-isotropic phase transition driven by the GS-cluster coupling. We also numerically compute and give heuristic insights into solution landscapes of confined BCN systems on 2D square domains, tailored by the GS-cluster coupling, temperature and external fields. This benchmark example illustrates the potential of this generalised model to capture tunable director profiles, cluster properties and potential biaxiality induced by antagonistic $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$-profiles.
@@ I Introduction
BCNs were discovered in the 1990’s; they are remarkable materials that can exhibit different types of order — polar, chiral, octupolar, biaxial order etc. They can exhibit structural transitions to more ordered phases —such as twist‑bend and smectic—depending on temperature and molecular architecture  [11]. These BCN phases offer tremendous potential for electro-optic devices, rheological applications and novel functional materials. Cybotactic clusters are an intrinsic feature of BCNs; they are smectic-like nanoclusters such that BCN molecules get locked into clusters because of their kinked shape. The existence of cybotactic clusters has been confirmed by multiple experimental methods, e.g., SAXS, cryo-TEM methods, electro-optic measurements [26, 24, 5] etc. Cybotactic clusters are small, usually tens of nanometers in lateral dimensions and persist in the nematic and sometimes, even in the isotropic BCN phase. Cybotactic clusters render additional ordering to the conventional nematic phase and this ordering clearly manifests in optical images (four-lobed structures, characteristic of local smectic‑C‑like ordering, in SAXS measurements), rheological properties (non-Newtonian properties), elastic properties (suppressed twist elastic constants leading to spontaneous chirality), macroscopic biaxiality [11] etc. As such, it is essential to quantify the additional ordering induced by the cybotactic clusters and how this ordering modulates the conventional ambient nematic ordering.
@@
In [17], Madhusudana proposes a two-state model for BCNs, based on the hypothesis that there are two types of molecules: the ground-state (GS) molecules that constitute the out-of-cluster ambient environment and the excited molecules (ES) that constitute the cybotactic clusters. There are two scalar order parameters, $S_{g}$ and $S_{c}$, to describe the orientational ordering of the GS molecules and the ES molecules respectively. Madhusudana assumes spatially uniform $S_{g}$ and $S_{c}$ profiles, i.e., he works with spatially homogeneous BCN systems along with various other assumptions — fixed time-independent and temperature-independent cluster sizes, an empirical coupling term between $S_{g}$ and $S_{c}$ and no explicit information about the BCN directors. Madhusudana proposes a phenomenological free energy comprising two Ginzburg-Landau type potentials for $S_{g}$ and $S_{c}$ with various temperature-dependent and material-dependent parameters and an empirical GS-ES coupling term that is parameterised by a coupling coefficient. The physically observable spatially homogeneous BCN configurations are modelled by minimisers of this two-state free energy. The most important finding concerns the existence of a new paranematic phase with small values of $S_{g}$ and $S_{c}$, just before the transition to the disordered isotropic phase. This model also effectively captures other experimental trends reported in [28] — enhanced order parameters induced by the cybotactic clusters, a nematic-paranematic phase transition followed by a paranematic-isotropic phase transition with increasing temperature and non-linear dependence of order parameters on external magnetic fields.
@@
However, Madhusudana’s model cannot account for confinement effects or spatial inhomogeneities. In [21], the authors propose a one-dimensional model for spatially inhomogeneous BCN samples, including elastic energy terms for $S_{g}$ and $S_{c}$. They study BCN samples in channel geometries with fixed boundary conditions for $S_{g}$ and $S_{c}$; their numerical results demonstrate that confinement or geometric frustration promotes interior clustering or enhances interior ordering and that the GS-ES coupling can promote ordered director profiles even at high temperatures contrary to the expectation of isotropy or disorder at high temperatures. In [19], the authors study the effects of doping on cybotactic clusters along with the emergence of polar order in such BCN systems. They perform extensive experiments on phase transitions, birefringence, order parameter measurements and dielectric spectroscopy data and propose a simple Madhusudana-type free energy to partially describe their experimental results. In this paper, we build on these works to propose a generalised free energy for confined BCN systems in two-dimensional (2D) and three-dimensional (3D) settings. By analogy with Landau-de Gennes theory, we define two tensor order parameters: $\mathbf{Q}_{g}$ to describe the orientational ordering of the GS molecules and $\mathbf{Q}_{c}$ to describe the additional ordering induced by the ES molecules within cybotactic clusters [18]. The key advantage is that $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ not only contain information about $S_{g}$ and $S_{c}$ but also contain information about the GS and ES directors, that are the preferred directions of average alignment of the GS molecules and new distinguished material directions induced by the cybotactic clusters. The free energy has an elastic energy to penalise spatial inhomogeneities and a bulk energy that determines the GS and ES ordering as a function of temperature, GS-ES coupling and other parameters, along with terms that account for external (electric or magnetic) fields. The bulk energy is a direct generalisation of Madhusudana’s free energy and is composed of Ginzburg-Landau type potentials for $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ respectively and a $\mathbf{Q}_{g}$-$\mathbf{Q}_{c}$ coupling term. The Ginzburg-Landau potentials dictate that $S_{g}$ and $S_{c}$ increase with decreasing temperatures and approach zero with increasing temperature (the isotropic or disordered phase). The GS-ES coupling and external field terms can strongly boost ordering, even for high temperatures. The physically observable textures are modelled by minimisers of this generalised energy, subject to the imposed boundary conditions. The minimisers and the critical points of the free energy are solutions of boundary-value problems for the associated Euler-Lagrange equations — typically a system of coupled and nonlinear partial differential equations and the rest of this manuscript is devoted to the study of critical points of this generalised free energy in physically motivated settings.
@@
The rest of the paper is organised as follows. In Section II, we introduce a generalised Landau-de Gennes type free energy for BCN systems. In Section III, we study phase transitions in spatially homogeneous BCN samples, with and without external fields, in 2D and 3D. Notably, we find two phase transitions in spatially homogeneous 3D BCN samples driven by increasing GS-ES coupling strength, at high temperatures. The GS-ES coupling strength is measured in terms of a parameter $\gamma$. If $\gamma=0$, then $S_{g}=S_{c}=0$ for spatially homogeneous samples at high temperatures (defined by $A>0$ in (5)). As $\gamma$ increases, the energy minimisers have positive values of $S_{g}$ and $S_{c}$ even for high temperatures. At a fixed high temperature, increasing $\gamma$ has a strong ordering effect that counteracts the disordering effects of high temperature. This is analogous to the nematic-paranematic phase transition and the paranematic-isotropic phase transition with increasing temperature, at fixed $\gamma$, as reported in [17]. The external fields typically boost the values of $S_{g}$ and $S_{c}$, shift the bifurcation points and we lose the paranematic-isotropic phase transition. Given that the values of $S_{g}$ and $S_{c}$ are very small in the paranematic phase, it is unclear if the paranematic phase would be distinguishable from the conventional isotropic phase in experiments. In Section IV, we study solution landscapes of confined BCN systems on a square domain with Dirichlet boundary conditions for $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$. This is a benchmark well-studied example for conventional nematics; see the papers [30, 16, 14]. A solution landscape is a network of critical points of the free‑energy surface, revealing possible pathways between the different stable solutions and unstable saddle points of the free energy. The unstable saddle points play a crucial role in the system dynamics. We study the interplay between domain size, temperature, the GS-ES coupling and external field strength on the energy minimisers and the associated solution landscapes. We recover some familiar $\mathbf{Q}_{g}$ profiles from the studies on conventional nematics, some new $\mathbf{Q}_{g}$ profiles tailored by the additional physics in this model, co-aligned $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ profiles with co-aligned GS and ES directors and antagonistic $(\mathbf{Q}_{g},\mathbf{Q}_{c})$-profiles with mutually perpendicular GS and ES directors. Our simple examples illustrate the tremendous potential of tuning model parameters to get exotic morphologies, some of which could lead to macroscopic biaxiality e.g., the antagonistic $(\mathbf{Q}_{g},\mathbf{Q}_{c})$ profiles. We conclude with a discussion of the limitations of the model and future improvements in Section VI.
@@ II Landau-de Gennes Modelling for Bent-Core Nematic Systems
In this paper, we propose a two-state generalised Landau-de Gennes (LdG) type free energy for confined BCN systems building on Madhusudana’s model in [17]. We define a GS LdG order parameter, $\mathbf{Q}_{g}$, which is a symmetric, traceless $d\times d$ matrix, where $d$ is the spatial dimension. The eigenvectors of $\mathbf{Q}_{g}$ model the GS nematic directors, interpreted as the locally preferred directions of averaged GS molecular alignment in space and the associated eigenvalues measure the degree of orientational ordering about the GS directors. We define the GS director to the eigenvector of $\mathbf{Q}_{g}$ with the largest positive eigenvalue. The GS is said to be in the isotropic phase if $\mathbf{Q}_{g}=0$. It is in the uniaxial phase if $\mathbf{Q}_{g}$ has two equal non-zero eigenvalues for $d=3$ and $\mathbf{Q}_{g}$ can be written as
@@
We treat the cybotactic clusters as uniformly distributed sub-nanometric/very small clusters, such that the size of the cluster is much smaller than the spacing between the clusters and the volume fraction of the clusters is much smaller than unity. This is the so-called “dilute limit” in homogenization theory [1], consistent with the assumptions in Madhusudana’s paper. In this case, the highly localised cybotactic clusters interact with the surrounding GS molecules by means of an averaged field, the $\mathbf{Q}_{c}$ - tensor field, referred to as the cluster order parameter in this paper. In principle, $\mathbf{Q}_{c}$ contains information about the cluster shapes and the GS-ES interactions, assuming that the clusters do not interact with each other. For example, in $d=2$, we have
@@
where $\mathbf{n}_{c}$ is referred to as the ES/cluster director and $S_{c}$ is positive. As an immediate consequence, we have two tensor order parameters, $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$, as opposed to two scalar order parameters, $S_{g}$ and $S_{c}$ in [17], that can describe complex information about the GS and cluster directors, structural transitions and confinement effects in inhomogeneous systems which are outside the scope of the two-state model in [17].
@@
where $\Omega$ is a $d$-dimensional bounded domain with Lipschitz boundary and $d=2,3$. The first two terms are the elastic energy density terms associated with $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ respectively, i.e., we assume a one-constant elastic energy density to penalise spatial inhomogeneities in $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ such as strong GS-ES interactions, interior defects, geometric frustration etc. $K_{g}$ and $K_{c}$ denote the positive elastic constants associated with $\mathbf{Q}_{g}$ and $\mathbf{Q}_{c}$ respectively. We note that other choices of anisotropic elastic energy densities are also possible, see [7]. For BCN materials, the bend elastic constant $K_{33}$ is often considerably smaller than the splay ($K_{11}$) or twist ($K_{22}$) constants, reflecting an intrinsic bend‑softening associated with the molecular geometry. The mole fraction of the ES molecules $a_{x}$ is given by
@@
as $E_{ex}$ is the excitation energy of the ES molecules, $k_{B}$ is Boltzmann’s constant and $T$ is the temperature. $a_{x}=0.1$ throughout the paper consistent with [21]. Further, there maybe an intrinsic $a_{x}$-dependence of the elastic constants. In the interest of brevity and accounting for uncertainties in the properties of the cybotactic clusters, including their elastic constants, we use the one-constant elastic energy density as a first approximation. The parameters $a_{g}$, $B_{g}$, $C_{g}$, and $T^{*}$ are the usual LdG parameters that describe the first-order nematic-isotropic transition for the GS molecules and $\gamma$ is the coupling parameter between the GS molecules and the clusters [21, 17]. We note that the term $\textrm{tr}\left(\mathbf{Q}_{g}\mathbf{Q}_{c}\right)$ is a generic coupling term for systems with two order parameters that is consistent with frame-indifference and material symmetry requirements [1]. $\alpha_{c}$ and $\beta_{c}$ are coefficients for the saturation terms to ensure that $|\mathbf{Q}_{c}|^{2}$ is reasonably bounded, for at least a certain range of $\gamma$. There is no evidence of first-order isotropic-nematic phase transitions within cybotactic clusters to date, so the term $tr\mathbf{Q}_{c}^{3}$ is absent. We also note that Madhusudana’s original work does not include a cubic term in $S_{c}$, analogous to a $tr\mathbf{Q}^{3}_{c}$ term. In the subsequent numerical simulations, we adjust the values of $\alpha_{c}$ and $\beta_{c}$ so that the resulting values of $S_{c}$ are reasonable and not too large. $N_{c}$ is the number of ES molecules within each cluster. We also consider the effects of an external field, $\mathbf{E}=E_{a}\mathbf{e}$, whose direction is modelled by a unit-vector $\mathbf{e}$, and the magnitude is $E_{a}$. $E_{el}$ is the external field energy density $(\frac{1}{2}\epsilon_{0}\Delta\epsilon E_{a}^{2})$ where $\epsilon_{0}$ is the free-space permittivity, $\Delta\epsilon$ is a material anisotropy parameter (dielectric or magnetic) for GS molecules and $J$ accounts for the larger shape anisotropy of ES molecules, i.e., $J>1$. In particular, the GS molecules tend to align with $\mathbf{E}$ if $\Delta\epsilon>0$ and orthogonal to $\mathbf{E}$ if $\Delta\epsilon<0$ [2].
@@
and $\tilde{\Omega}$ is the rescaled domain. For comparison with previous work on one-dimensional BCN systems in [21], we set $A=\pm 0.04$, $B=0.34$, $C=0.9$, $D=2\times 10^{-3}$, $E=1.6\times 10^{-4}$, $M=9.7\times 10^{-5}$, $N=1.78\times 10^{-3}$, $P=0.133$ which are derived from $K_{g}=K_{c}=K=15\mathrm{pN}=15\times 10^{-12}\mathrm{N}=15\times 10^{-7}\mathrm{dyn}\text{(under one constant approximation)}$; $a_{g}=0.04\times 10^{7}/4\ \mathrm{ergs/(cm^{3}K)}$, $B_{g}=1.7\times 10^{7}/4\ \mathrm{ergs/cm^{3}}$, $C_{g}=4.5\times 10^{7}/4\ \mathrm{ergs/cm^{3}}$, $\alpha_{c}=0.22\times 10^{7}/4\ \mathrm{ergs/cm^{3}}$, $\beta_{c}=4.0\times 10^{7}/4\ \mathrm{ergs/cm^{3}}$, and $\gamma=5.0\times 10^{7}/4\ \mathrm{ergs/cm^{3}}$; $T^{*}=355K$, $N_{c}=50$, $J=1.2$, $E_{el}=2000\mathrm{ergs/cm^{3}}$, $E_{ex}=1.1\times 10^{-13}\mathrm{ergs}$, and $T=360\mathrm{K}$ (for the high temperature case with $A>0$ ) and $350\mathrm{K}$ (for the low temperature case with $A<0$). In subsequent sections, we focus on phase transitions in spatially homogeneous systems and confinement effects in square domains as a function of temperature $A$, the GS-ES coupling strength measured by $D$ and external field strength, $E$.
For clarity, we temporarily refer to the $F$ from Eq. (4), where $\Omega$ is a $d$-dimensional region, as $F_{d}$. In three dimensions, the free energy $F_{3}$ has the unit of energy (erg), whereas in two dimensions, $F_{2}$ has the unit of force (N). This effectively assumes a uniform extension along the third dimension, so the true three‑dimensional energy can be written as $F_{3}=F_{2}\times(\text{thickness})$. After nondimensionalization in Eq. (5), however, the factor $\lambda^{2-d}$, which depends on the spatial dimension $d$, renders the dimensionless free energy $\tilde{F}$ unit‑free in both two and three dimensions.